OpenStudy (anonymous):

Consider the function represented by the equation 6q = 3s - 9. Write the equation in function notation, where q is the independent variable. f(q) = 2s + 3 f(q) = 2q + 3

OpenStudy (anonymous):

f(q)=1/2q-3/2 f(s)=1/2s-3/2

OpenStudy (anonymous):

@Blonde_Gangsta

OpenStudy (anonymous):

@Godlovesme @JackofallTradez

OpenStudy (jackofalltradez):

I do believe the correct answer is D, f(q) = 1/2s - 3/2

OpenStudy (anonymous):

thanks can u help me with 2 more ?

OpenStudy (jackofalltradez):

Sure ^_^

OpenStudy (anonymous):

Consider the function represented by the equation . Which answer shows the equation written in function notation with x as the independent variable?

OpenStudy (anonymous):

f(x)=6x+9 f(x)=1/6x+3/2 f(y)=6y+9 f(y)=1/6y+3/2

OpenStudy (anonymous):

The function V(r)=4/3piR^3 can be used to find the volume of air inside a basketball given its radius. What does V(r) represent? the radius of the basketball when the volume is V the volume of the basketball when the radius is r the volume of the basketball when the radius is V the radius of the basketball when the volume is r

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@DenishaH

OpenStudy (jackofalltradez):

The function V(r)=4/3piR^3 can be used to find the volume of air inside a basketball given its radius. What does V(r) represent? B.) the volume of the basketball when the radius is r

OpenStudy (jackofalltradez):

This is because V(r) means 'the volume of [Object] when the the radius is r

OpenStudy (anonymous):

thanks a lot :) @JackofallTradez